How do you solve this system of equations:
y=3x+1
x+2y=-5


Answer by rfer(16322)

y=3x+1
x+2y=-5
x+2(3x+1)=-5
x+6x+2=-5
7x=-7
x=-1
-------------
-1+2y=-5
2y=-4
y=-2
(-1,-2)
------------
y=3x+1
-2=3*-1+1
-2=-2